What are the units for the corresponding reactions: zero-order, first-order, second-order, and third-order reaction?


1
Expert's answer
2021-03-25T04:09:02-0400

Zero order constant = Ms-1

First order constant= s-1

Second order constant = M-1s-1

Third order constant = M-2s-1